# How Useful is Continued Pre-Training for Generative Unsupervised Domain Adaptation?

### v2

@@ Abstract
Recent breakthroughs in scale have enabled the emergence of powerful generative language models, and the ability to fine-tune these models on various tasks by casting them into prompts or instructions. In this landscape, the problem of Unsupervised Domain Adaptation (UDA), or the problem of leveraging knowledge from a labeled source domain to an unlabeled target domain, has been left behind, with recent UDA methods still addressing discriminative classification. In particular, two popular UDA approaches, involving Continued Pre-Training (CPT) and learning domain invariant representations, have been under-explored in the generative setting, signaling a gap. In this work, we evaluate the utility of CPT for generative UDA. We first perform an empirical evaluation to measure the trade-offs between CPT and strong methods promoting domain invariance. We further evaluate how well the benefits of CPT extend to different architectures, tuning methods and data regimes. We then motivate the use of CPT by studying to what degree it benefits classification performance on the target domain. Finally, we attempt to understand the mechanism behind which CPT improves classification performance on the unlabeled target domain. Our findings suggest that a implicitly learns the downstream task while predicting masked words informative to that task. Our work connects the body of UDA research with that of instruction tuning, enabling an initial step towards a wider applicability of modern language models.
@@ 1 Introduction
Recent advancements in the pre-training of language models have enabled the widespread use of powerful generative models, which can be leveraged across multiple domains with no training (Brown et al., 2020; Scao et al., 2022; Touvron et al., 2023, inter alia). Despite these advancements, these autoregressive models are still fragile under certain kinds of data distribution shifts, making their applications across these domains challenging (Ribeiro et al., 2020). This is addressed, in part, by the concept of instruction tuning with templates (Zhang et al., 2023; Sanh et al., 2022; Ouyang et al., 2022; Wang et al., 2022; Wei et al., 2022), enabling the learning of new tasks without any randomly initialized parameters.
@@
The problem of unsupervised domain adaptation (UDA) leverages learned knowledge from a labeled source domain to an unlabeled target domain (Pan & Yang, 2010; Ganin & Lempitsky, 2015; Long et al., 2015, inter alia). It is useful for adaptation to unlabeled domains with high labeling costs, where supervised instruction tuning does not suffice. Despite the pervasive need for models to generalize to such domains, recent UDA methods still address discriminative classification, barring the application of these approaches to recent generative models. In particular, Continued Pre-Training and Domain Invariance based methods, two widely popular classes of UDA approaches (Ramponi & Plank, 2020), are completely unexplored in the generative setting.
@@
The Continued Pre-Training (CPT) approach involves extended pre-training on a domain or task, followed by supervised training on the downstream task (Gururangan et al., 2020). This approach has been widely used for adaptation to labeled domains (Gao et al., 2021; Kim et al., 2021; Hung et al., 2023, inter alia), and in the UDA setup for unlabeled domains (Han & Eisenstein, 2019; Zhang et al., 2021b; Karouzos et al., 2021; Pfeiffer et al., 2020; Parović et al., 2023). Invariance based approaches attempt to learn representations that are invariant across domains (Tzeng et al., 2014; Ganin et al., 2016; Wu & Shi, 2022; Guo et al., 2022), with the notion that when the learned representations from both domains cannot be distinguished by a classifier and the classifier performs well on the source domain, it will also exhibit strong performance on the target domain. These two classes of methods introduce a trade-off: invariance based methods suffer from instability issues (Han & Eisenstein, 2019; Kashyap et al., 2021), while continued pre-training requires a larger computational budget. But how would this trade-off extrapolate to the generative setting? For example, invariance based methods are well motivated in discriminative tasks, where there is a clear decision boundary; however, the same does not hold for generative tasks.
@@
To address these gaps, we introduce the setting of Generative UDA, where an autoregressive model is trained to leverage knowledge from a labeled source domain to an unlabeled target domain, using only next word prediction as its objective. We formalize the use of CPT for this setting in Section 2, and then attempt to explore and understand the behaviour of CPT for Generative UDA. We begin by performing an empirical analysis on 40 real-world domain pairs to explore the tradeoff between continued pre-training and invariance based approaches, and find vanilla CPT to be competitive with and significantly more stable than a state of the art invariance based approach (Section 3). We then stress test CPT, by applying it across varying model architectures and scales, tuning approaches and data regimes. We find that CPT is robust across these settings, unlike our invariance based approach (Section 4).
@@ 2.2 CPT for UDA as a Sequence of Prompt Based Tasks
We reuse the input sequences from the source-labeled dataset $\mathcal{D}_{\text{src}}$ as the source-unlabeled dataset, denoted as $\mathcal{D}_{\text{src}}^{\mathcal{X}}$. Next, similar to Raffel et al. (2020); Song et al. (2019), for an unlabeled sequence $x\in\mathcal{D}_{\text{src}}^{\mathcal{X}}$ and $\mathcal{D}_{\text{src}}$, we use a prompt template to convert the sequence $x$ to an input-output sequence pair, i.e., $\mathbb{M}(x)=(\tilde{x},\tilde{y})$. For masked language modeling (MLM), an instruction is pre-pended to a randomly masked sequence $x$ to create $\tilde{x}$. The output sequence $\tilde{y}$ is a concatenation of masked words from $x$. For example, given $x=$“The movie was so cool! Two hours of fun.”, we construct $\tilde{x}=$“Fill in the blanks: ”The _ cool! Two hours _”, and $\tilde{y}=$ <sep> movie was so <sep> of fun. <sep>. For causal language modeling (CLM), $\tilde{x}=x$.
@@
$\ell(\tilde{x},\tilde{y};\theta)=-\frac{1}{|\tilde{y}|}\sum_{t}\log P_{\theta}(\tilde{y}_{t}|\tilde{x},\tilde{y}_{1:t-1});~{}~{}~{}~{}~{}~{}\mathcal{L}_{\text{CPT}}(\mathcal{D};\theta)={\frac{1}{|\mathcal{D}|}}\sum_{x\in\mathcal{D}}\ell(\mathbb{M}(x);\theta).$ (1)
@@
$\mathcal{L}_{\text{CLS}}(\mathcal{D}_{\text{src}};\theta)={\frac{1}{|\mathcal{D}_{\text{src}}|}}\sum_{(x,y)\in\mathcal{D}_{\text{src}}}l(\mathbb{C}(x,y);\theta).$ (2)
@@ 3.1 Experimental Setup
Since our goal is to study the behaviour of CPT for generative UDA, we it compare with a simple supervised baseline, and a state-of-the-art invariance based approach.
@@
  • UDAPTER: Malik et al. (2023) propose an invariance based method that measures the multi-kernel maximum mean discrepancy (MK-MMD) (Gretton et al., 2012; Bousmalis et al., 2016) between source and target embeddings from each transformer layer and sums them to obtain an aggregate loss $\mathcal{L}_{\text{div}}$. The final loss is the weighted sum of $\mathcal{L}_{\text{div}}$ and the classification loss, i.e., $\mathcal{L}=\lambda\;\mathcal{L}_{\text{cls}}+(1-\lambda)\;\mathcal{L}_{\text{div}}$, where $\lambda$ gradually changes from 0 to 1 during training. Here, we use the embeddings from a model as it is being instruction tuned on the downstream classification task. Their method achieves state-of-the-art performance, and outperforms popular UDA approaches like DANN (Ganin et al., 2016) and DSN (Bousmalis et al., 2016); thus we only compare CPT with this approach.
@@ 3.2 Continued Pre-training is Competitive with Domain-Invariance Methods
Table 1: Avg. target-domain classification accuracy and standard deviation over 3 runs on Amazon Review and MNLI.
Performance
@@
CPT performs more stably than UDAPTER, with the invariance-based method often reporting a variance of over 20% across runs (Table 6 in Appendix C). For example, for the MNLI pair Fiction (F) $\to$ Government (G), minimizing UDAPTER yields a variance of 23.4% across runs. This observation is consistent with existing findings (Kashyap et al., 2021; Han & Eisenstein, 2019) that minimizing divergence measures like MMD, when combined with auxiliary task-specific loss functions, result in training instabilities and vanishing gradients. We discuss this in more detail in Appendix J.
@@ 4 How General are the Benefits of CPT?
We extend our analysis from MLM with encoder-decoder language models, to causal language modeling (CLM) with decoder-only language models, using GPT-2 (medium) (Radford et al., 2019). We perform CLM in the first training phase by simply training the model for next-word prediction given the original sequence. Table 2 shows that CPT provides strong improvements on the target domain in comparison to the invariance based baseline.
@@
CPT Benefits are Consistent across Tuning Approaches.
PEFT approaches have been shown to introduce resilience to domain shift (Fu et al., 2023). To isolate this effect from the CPT framework, we use T5v1.1 to evaluate CPT in a full fine-tuning setup. Additionally, we compare CPT with two PEFT approaches [^4] We choose Adapters because He et al. (2022) present a unified view of PEFT approaches which shows that the operations applied by Adapters are very similar to those of Prefix Tuning (Li & Liang, 2021) and LoRA (Hu et al., 2022). We choose (IA)3 since it is a state-of-the-art PEFT approach that uses a fraction of the learnable parameters of Adapters (More in Appendix E). : Adapters (Houlsby et al., 2019) and (IA)3 (Liu et al., 2022). We see in Figure 2 (Table 9 in Appendix E) that CPT continues to perform stronger than the domain invariance based UDAPTER method.
@@ 6 Why does Word Prediction Aid Classification for Generative UDA?
To test this hypothesis, we quantize the “informativeness” of each word to a classification task: an informative word is highly correlated with any of the labels in the downstream task. [^6] These informative words are similar to pivot features (Blitzer et al., 2006; Ziser & Reichart, 2018; Ben-David et al., 2020, inter alia), with the exception that they are chosen based on information from the source domain only. Specifically, we follow Gururangan et al. (2018) and use pointwise mutual information (PMI) (Fano, 1961) of the word with respect to the class label:
@@ 7 Discussion
Our results in Section 3.2 show that continued pre-training and methods promoting domain invariance are competitive with each other. Continued pre-training suffers from the computational drawback of requiring an additional phase of training. Conversely, invariance based methods are difficult to optimize, possibly requiring more runs to achieve a stable optimum, and having a higher amortized computational cost. Inspired by Karouzos et al. (2021), we introduce a simple single phase variant of continued pre-training which is equivalent in performance to its two phase variant, nullifying the additional computational overhead of the approach (more details in Appendix I).
@@
Recent breakthroughs in scale have showcased that large language models (LLMs) are highly powerful, and can perform various downstream tasks with limited or no training. This may raise question on the relevance of the UDA problem as a whole — does a model even require expensive adaptation to a domain it may have already been exposed to during pre-training? In addition to our analysis in Section 5, we argue that the requirement to adapt small models to unseen domains still holds in specific cases. Small supervised models have been shown to outperform zero-shot general-purpose LLMs on various downstream tasks (Huang et al., 2023; Zhu et al., 2023), and safety critical domains like healthcare and finance would benefit more from these models than a generalist LLM. Our study does not address how to better adapt to domains, rather we investigate ways a model may adapt to data unseen during pre-training. This is a question that holds for current LLMs, and will continue to hold as long as models are unable to access infinite data during pre-training.
@@ 8 Related Work
A major class of approaches in Model-centric UDA methods (Ramponi & Plank, 2020) aims to minimize $\mathcal{H}\Delta\mathcal{H}$ divergence (Ben-David et al., 2010) between the source and target domain features, through adversarial training (Tzeng et al., 2014; Ganin et al., 2016; Tzeng et al., 2017; Guo et al., 2022, inter alia) or through minimizing measures of domain similarity (Bousmalis et al., 2016; Ge et al., 2023). Malik et al. (2023) have shown the minimization of MMD to outperform other invariance based methods. However, past work has shown that domain-invariance is a weak constraint for adaptation (Zhao et al., 2019; Karouzos et al., 2021), could introduce domain-specific hyperparameters (Trung et al., 2022), and is also prone to instability issues (Han & Eisenstein, 2019; Sun et al., 2019; Wilson & Cook, 2020; Kashyap et al., 2021).
@@
The limitations of invariance-based model-centric methods have encouraged the emergence of alternate approaches, based on self-supervised learning through contrastive learning (Kumar et al., 2022; Shen et al., 2022; Long et al., 2022), pseudo-labeling (Zhou & Li, 2005; Ruder & Plank, 2017, inter alia) or language model pre-training. Despite not being directly useful to certain downstream tasks (Uppaal et al., 2023), CPT has been used for adaptation to labeled tasks, in both full fine-tuning (Gururangan et al., 2020; Lee et al., 2020; Gao et al., 2021, inter alia) and PEFT setups (Kim et al., 2021; Hung et al., 2023). A smaller body of work has explored the utility of CPT in a UDA setup (Han & Eisenstein, 2019; Zhang et al., 2021b; Karouzos et al., 2021; Parović et al., 2023), identifying the class of methods to be more stable than invariance-based methods.

### v3

@@ Abstract
Recent breakthroughs in scale have enabled the emergence of powerful generative language models, and the ability to fine-tune these models on various tasks by casting them into prompts or instructions. In this landscape, the problem of Unsupervised Domain Adaptation (UDA), or the problem of leveraging knowledge from a labeled source domain to an unlabeled target domain, has been left behind, with recent UDA methods still addressing discriminative classification. In particular, two popular UDA approaches, involving Continued Pre-Training (CPT) and learning domain invariant representations, have been under-explored in the generative setting, signaling a gap. In this work, we evaluate the utility of CPT for generative UDA. We first perform an empirical evaluation to measure the trade-offs between CPT and strong methods promoting domain invariance. We further evaluate how well the benefits of CPT extend to different architectures, tuning methods and data regimes. We then motivate the use of CPT by studying to what degree it benefits classification performance on the target domain. Finally, we attempt to understand the mechanism behind which CPT improves classification performance on the unlabeled target domain. Our findings suggest that the model implicitly learns the downstream task while predicting masked words informative to that task. Our work connects the body of UDA research with that of instruction tuning, enabling an initial step towards a wider applicability of modern language models. Our code is available at https://github.com/Uppaal/cpt-generative-uda.
@@ 1 Introduction
Recent advancements in the pre-training of language models have enabled the widespread use of powerful generative models, which can be leveraged across multiple domains with no training (Brown et al., 2020; Scao et al., 2022; Touvron et al., 2023, inter alia). Despite these advancements, these autoregressive models are still fragile under certain kinds of data distribution shifts, making their applications across these domains challenging (Ribeiro et al., 2020; Bajaj et al., 2021; Chuang et al., 2023; Uppaal et al., 2024, inter alia). This is addressed, in part, by the concept of instruction tuning with templates (Zhang et al., 2023; Sanh et al., 2022; Ouyang et al., 2022; Wang et al., 2022; Wei et al., 2022), enabling the learning of new tasks without any randomly initialized parameters.
@@
The problem of unsupervised domain adaptation (UDA) leverages learned knowledge from a labeled source domain to an unlabeled target domain (Pan and Yang, 2010; Ganin and Lempitsky, 2015; Long et al., 2015, inter alia). It is useful for adaptation to unlabeled domains with high labeling costs, where supervised instruction tuning does not suffice. Despite the pervasive need for models to generalize to such domains, recent UDA methods still address discriminative classification, barring the application of these approaches to recent generative models. In particular, Continued Pre-Training and Domain Invariance-based methods, two widely popular classes of UDA approaches (Ramponi and Plank, 2020), are completely unexplored for UDA in the generative setting.
@@
The Continued Pre-Training (CPT) approach involves extended pre-training on a domain or task, followed by supervised training on the downstream task (Gururangan et al., 2020). This approach has been widely used for adaptation to labeled domains (Gao et al., 2021; Kim et al., 2021; Hung et al., 2023, inter alia), and in the UDA setup for unlabeled domains (Han and Eisenstein, 2019; Zhang et al., 2021b; Karouzos et al., 2021; Pfeiffer et al., 2020; Parović et al., 2023). Invariance-based approaches attempt to learn representations that are invariant across domains (Tzeng et al., 2014; Ganin et al., 2016; Wu and Shi, 2022; Guo et al., 2022), with the notion that when the learned representations from both domains cannot be distinguished by a classifier and the classifier performs well on the source domain, it will also exhibit strong performance on the target domain. These two classes of methods introduce a trade-off: invariance-based methods suffer from instability issues (Han and Eisenstein, 2019; Kashyap et al., 2021), while continued pre-training requires a larger computational budget. But how would this trade-off extrapolate to the generative setting? For example, invariance-based methods are well motivated in discriminative tasks, where there is a clear decision boundary; however, the same does not hold for generative tasks.
@@
To address these gaps, we introduce the setting of Generative UDA, where an autoregressive model is trained to leverage knowledge from a labeled source domain to an unlabeled target domain, using only next word prediction as its objective. We formalize the use of CPT for this setting in Section 2, and then attempt to explore and understand the behaviour of CPT for Generative UDA. We begin by performing an empirical analysis on 40 real-world domain pairs to explore the tradeoff between continued pre-training and invariance-based approaches, and find vanilla CPT to be competitive with and significantly more stable than a state of the art invariance-based approach (Section 3). We then stress test CPT, by applying it across varying model architectures and scales, tuning approaches and data regimes. We find that CPT is robust across these settings, unlike our invariance-based approach (Section 4).
@@ 2.2 CPT for UDA as a Sequence of Prompt Based Tasks
We reuse the input sequences from the source-labeled dataset $\mathcal{D}_{\text{src}}$ as the source-unlabeled dataset, denoted as $\mathcal{D}_{\text{src}}^{\mathcal{X}}$. Next, similar to Raffel et al. (2020); Song et al. (2019), for an unlabeled sequence $x\in\mathcal{D}_{\text{src}}^{\mathcal{X}}$ and $\mathcal{D}_{\text{src}}$, we use a prompt template to convert the sequence $x$ to an input-output sequence pair, i.e., $\mathbb{M}(x)=(\tilde{x},\tilde{y})$. For masked language modeling (MLM), an instruction is pre-pended to a randomly masked sequence $x$ to create $\tilde{x}$. The output sequence $\tilde{y}$ is a concatenation of masked words from $x$. For example, given $x=$“The movie was so cool! Two hours of fun.”, we construct $\tilde{x}=$“Fill in the blanks: "The _ cool! Two hours _”, and $\tilde{y}=$ <sep> movie was so <sep> of fun. <sep>. For causal language modeling (CLM), $\tilde{x}=x$.
@@
$\ell(\tilde{x},\tilde{y};\theta)=-\frac{1}{|\tilde{y}|}\sum_{t}\log P_{\theta}(\tilde{y}_{t}|\tilde{x},\tilde{y}_{1:t-1})$ (1)
$\mathcal{L}_{\text{CPT}}(\mathcal{D};\theta)={\frac{1}{|\mathcal{D}|}}\sum_{x\in\mathcal{D}}\ell(\mathbb{M}(x);\theta)$
@@
$\mathcal{L}_{\text{CLS}}(\mathcal{D}_{\text{src}};\theta)={\frac{1}{|\mathcal{D}_{\text{src}}|}}\sum_{(x,y)\in\mathcal{D}_{\text{src}}}l(\mathbb{C}(x,y);\theta)$ (2)
@@ 3.1 Experimental Setup
Since our goal is to study the behaviour of CPT for generative UDA, we compare it with a simple supervised baseline, and a state-of-the-art invariance-based approach.
@@
  • UDAPTER: Malik et al. (2023) propose an invariance-based method that measures the multi-kernel maximum mean discrepancy (MK-MMD) (Gretton et al., 2012; Bousmalis et al., 2016) between source and target embeddings from each transformer layer and sums them to obtain an aggregate loss $\mathcal{L}_{\text{div}}$. The final loss is the weighted sum of $\mathcal{L}_{\text{div}}$ and the classification loss, i.e., $\mathcal{L}=\lambda\;\mathcal{L}_{\text{cls}}+(1-\lambda)\;\mathcal{L}_{\text{div}}$, where $\lambda$ gradually changes from 0 to 1 during training. Here, we use the embeddings from a model as it is being instruction tuned on the downstream classification task. Their method achieves state-of-the-art performance, and outperforms popular UDA approaches like DANN (Ganin et al., 2016) and DSN (Bousmalis et al., 2016); thus we only compare CPT with this approach.
@@ 3.2 Continued Pre-training is Competitive with Domain-Invariance Methods
Table 1: Avg. target-domain classification accuracy and standard deviation over 3 runs.
@@
CPT performs more stably than UDAPTER, with the invariance-based method often reporting a variance of over 20% across runs (Table 6 in Appendix C). For example, for the MNLI pair Fiction (F) $\to$ Government (G), minimizing UDAPTER yields a variance of 23.4% across runs. This observation is consistent with existing findings (Kashyap et al., 2021; Han and Eisenstein, 2019) that minimizing divergence measures like MMD, when combined with auxiliary task-specific loss functions, result in training instabilities and vanishing gradients. We discuss this in more detail in Appendix J.
@@ 4 How General are the Benefits of CPT?
We extend our analysis from MLM with encoder-decoder language models, to causal language modeling (CLM) with decoder-only language models, using GPT-2 (medium) (Radford et al., 2019). We perform CLM in the first training phase by simply training the model for next-word prediction given the original sequence. Table 2 shows that CPT provides strong improvements on the target domain in comparison to the invariance-based baseline.
@@
CPT Benefits are Consistent across Tuning Approaches.
PEFT approaches have been shown to introduce resilience to domain shift (Fu et al., 2023). To isolate this effect from the CPT framework, we use T5v1.1 to evaluate CPT in a full fine-tuning setup. Additionally, we compare CPT with two PEFT approaches [^4] We choose Adapters because He et al. (2022) present a unified view of PEFT approaches which shows that the operations applied by Adapters are very similar to those of Prefix Tuning (Li and Liang, 2021) and LoRA (Hu et al., 2022). We choose (IA)3 since it is a state-of-the-art PEFT approach that uses a fraction of the learnable parameters of Adapters (More in Appendix E). : Adapters (Houlsby et al., 2019) and (IA)3 (Liu et al., 2022). We see in Figure 2 (Table 9 in Appendix E) that CPT continues to perform stronger than the domain invariance-based UDAPTER method.
@@ 6 Why does Word Prediction Aid Classification for Generative UDA?
To test this hypothesis, we quantize the “informativeness” of each word to a classification task: an informative word is highly correlated with any of the labels in the downstream task. [^6] These informative words are similar to pivot features (Blitzer et al., 2006; Ziser and Reichart, 2018; Ben-David et al., 2020, inter alia), with the exception that they are chosen based on information from the source domain only. Specifically, we follow Gururangan et al. (2018) and use pointwise mutual information (PMI) (Fano, 1961) of the word with respect to the class label:
@@ 7 Discussion
Our results in Section 3.2 show that continued pre-training and methods promoting domain invariance are competitive with each other. Continued pre-training suffers from the computational drawback of requiring an additional phase of training. Conversely, invariance-based methods are difficult to optimize, possibly requiring more runs to achieve a stable optimum, and having a higher amortized computational cost. Inspired by Karouzos et al. (2021), we introduce a simple single phase variant of continued pre-training which is equivalent in performance to its two phase variant, nullifying the additional computational overhead of the approach (more details in Appendix I).
@@
Recent breakthroughs in scale have showcased that large language models (LLMs) are highly powerful, and can perform various downstream tasks with limited or no training. This may raise question on the relevance of the UDA problem as a whole — does a model even require expensive adaptation to a domain it may have already been exposed to during pre-training? In addition to our analysis in Section 5, we argue that the requirement to adapt small models to unseen domains still holds in specific cases. Small supervised models have been shown to be comparable with, or even outperform, zero-shot general-purpose LLMs on various downstream tasks (Huang et al., 2023; Zhu et al., 2023; Tang et al., 2024), serving as lightweight and customizeable (through fine-tuning) alternatives. Safety critical domains like healthcare and finance would benefit more from these models than a generalist LLM. Our study does not address how to better adapt to domains, rather we investigate ways a model may adapt to data unseen during pre-training. This is a question that holds for current LLMs, and will continue to hold as long as models are unable to access infinite data during pre-training.
@@ 8 Related Work
A major class of approaches in Model-centric UDA methods (Ramponi and Plank, 2020) aims to minimize $\mathcal{H}\Delta\mathcal{H}$ divergence (Ben-David et al., 2010) between the source and target domain features, through adversarial training (Tzeng et al., 2014; Ganin et al., 2016; Tzeng et al., 2017; Guo et al., 2022, inter alia) or through minimizing measures of domain similarity (Bousmalis et al., 2016; Ge et al., 2023). Malik et al. (2023) have shown the minimization of MMD to outperform other invariance-based methods. However, past work has shown that domain-invariance is a weak constraint for adaptation (Zhao et al., 2019; Karouzos et al., 2021), could introduce domain-specific hyperparameters (Trung et al., 2022), and is also prone to instability issues (Han and Eisenstein, 2019; Sun et al., 2019; Wilson and Cook, 2020; Kashyap et al., 2021).
@@
The limitations of invariance-based model-centric methods have encouraged the emergence of alternate approaches, based on self-supervised learning through contrastive learning (Kumar et al., 2022; Shen et al., 2022; Long et al., 2022), pseudo-labeling (Zhou and Li, 2005; Ruder and Plank, 2017, inter alia) or language model pre-training. Despite not being directly useful to certain downstream tasks (Uppaal et al., 2023), CPT has been used for adaptation to labeled tasks, in both full fine-tuning (Gururangan et al., 2020; Lee et al., 2020; Gao et al., 2021, inter alia) and PEFT setups (Kim et al., 2021; Hung et al., 2023). A smaller body of work has explored the utility of CPT in a UDA setup (Han and Eisenstein, 2019; Zhang et al., 2021b; Karouzos et al., 2021; Parović et al., 2023), identifying the class of methods to be more stable than invariance-based methods.